OpenStudy (anonymous):

Solve 20x-18<50<36x-16

OpenStudy (anonymous):

Alright. Treat this inequality as an equation. Split it up piece by piece. 20x-18=50 So, subtract 50 from both sides first. :) What do you get?

OpenStudy (anonymous):

i got -32

OpenStudy (anonymous):

Don't solve, just subtract 50 from both sides. :)

OpenStudy (anonymous):

Oh wait, is that your final answer?

OpenStudy (anonymous):

yes i guess

hero (hero):

Solve both: 20x -18 < 50 36x -16 > 50

OpenStudy (anonymous):

ok now wat

hero (hero):

Did you solve both yet? If so, what did you get?

OpenStudy (anonymous):

ok 20-36= -16 and 18-16= 2 and 50- 50=0 right am i on the right track

hero (hero):

That's simply not the right approach. Solve this first: 20x - 18 < 50

hero (hero):

Then after solving that, solve this: 36x - 16 > 50

hero (hero):

Do not change any of the signs to equal signs while solving